OpenStudy (anonymous):

Please help with this math question!! The volume of a cone is 300 cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use 3.14 for π. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

its 1000000000000000000

OpenStudy (anonymous):

volume of a cone is =1/3(πr^2h) 300=1/3*3.14*9*9*h h=3*300/3.14*9*9 h=3.53 km

OpenStudy (anonymous):

Thank you! :)